OpenStudy (anonymous):

Please Help c: What are the solutions to the following system of equations? y = x2 + 12x + 30 8x − y = 10 (−4, −2) and (2, 5) (−2, −4) and (2, 5) (−2, −4) and (5, 2) No Real Solutions

OpenStudy (anonymous):

What is the solution to the following system of equations? y = x2 + 10x + 11 y = x2 + x − 7 (−2, −5) (2, −5) (−2, 5)

OpenStudy (campbell_st):

rewrite the 2nd equation with y as the subject y = 8x - 10 now equate it to the 1st equation and you get \[8x - 10 = x^2 + 12x + 10\] now you have 1 equation in 1 unknown to solve

OpenStudy (anonymous):

y = x2 + 12x + 30 8x − y = 10 express y from second equation: \(y=8x-10\) substitute in the 1º equation: \(8x-10=x^2 + 12x + 30\) rearange terms: \(x^2+4x+40=0\) solve this. You will get 2 solutions, which later substitute in the 2º equation to get corresponding y value

OpenStudy (campbell_st):

oops should be ' \[8x - 10 = x^2 + 12x + 30\] just solve for x

OpenStudy (anonymous):

Mm.. ill try. One sec, ill let you know what I get

OpenStudy (missmob):

:O maths

OpenStudy (campbell_st):

and can you check your equations as they may be a typo in them...

OpenStudy (nirmalnema):

you can get the ans by by making the quadratic eq..in x

OpenStudy (missmob):

we will put all the x on one side

OpenStudy (anonymous):

The what? LOL

OpenStudy (campbell_st):

because based on the information there are any solutions that are real numbers

OpenStudy (campbell_st):

oops... which is an option

OpenStudy (campbell_st):

sorry...

OpenStudy (anonymous):

Ok.. so its D.. cool

OpenStudy (missmob):

maths is confusing

OpenStudy (anonymous):

But idk how to get D

OpenStudy (missmob):

*takes out the gun shoots maths*

OpenStudy (campbell_st):

well do you know about the discriminant...? thats how you check...

OpenStudy (anonymous):

lol...

OpenStudy (anonymous):

No.. but I got 8x-10=x2+12x+30 so far..

OpenStudy (anonymous):

What do I do from there?

OpenStudy (campbell_st):

subtract 8x from both sides... then add 10 to both sides of the equation...

OpenStudy (anonymous):

Oh.. I knew that. For some reason I thought I needed to bring my X from the right to the left.

OpenStudy (anonymous):

Ok.. so I got x=x2+4x+40 Is that right?

OpenStudy (nirmalnema):

ya you got it right

OpenStudy (campbell_st):

and the discriminant is for a quadratic \[ax^2 + bx + c = 0\] thew discriminant is \[b^2 - 4ac\] you have a = 1, b = 4 and c = 40 it will be less than zero... meaning complex roots

OpenStudy (anonymous):

Sooo its D.

OpenStudy (anonymous):

Thanks everyone! c: